OpenStudy (anonymous):

Verify: sin(2π - α) = -sinα

OpenStudy (anonymous):

Do you still need help?

OpenStudy (anonymous):

Yup, still new to these.

OpenStudy (anonymous):

ok

OpenStudy (calculusfunctions):

Expand the left side by using the identity sin (A + B) = (sin A)(cos B) + (cos A)(sin B)

OpenStudy (calculusfunctions):

@workin_daily ? Do you understand?

OpenStudy (anonymous):

Yeah I think I can do it now that I know my starting point.

OpenStudy (calculusfunctions):

Good, Go ahead! Let me know if you need further assistance.

OpenStudy (anonymous):

But shouldn't the plus signs be minus for this problem (in the identity formula for sin)?

OpenStudy (anonymous):

Just checking

OpenStudy (calculusfunctions):

Nope! What I gave you is correct!

OpenStudy (anonymous):

Wait.. So basicly 2π is beta, and it's (Beta + (-alpha)) ?

OpenStudy (anonymous):

And that's why the formula keeps the plus?

OpenStudy (calculusfunctions):

For your question sin (A - B) = (sin A)(cos B) - (cos A)(sin B) You just replace the plus sign with a minus sign.

OpenStudy (calculusfunctions):

OH I just realized, that's what you meant earlier when you asked "shouldn't the plus signs be minus for this problem (in the identity formula for sin)?"

OpenStudy (anonymous):

That's what I was wondering at first but since the problem started out with it minusing alpha and not beta, if that means it's actually beta + negative alpha

OpenStudy (calculusfunctions):

Yes Yes! Sorry, am a bit tired.

OpenStudy (anonymous):

Same here ha ha.. so: \[\sin(2\pi+(-\alpha))=\sin(\alpha)\cos(\beta)+\cos(\alpha)\sin(\beta)\] beta in the formula being 2π

OpenStudy (calculusfunctions):

You're making it too complicated. sin (A - B) = (sin A)(cos B) - (cos A)(sin B) Just use this directly to obtain\[\sin (2\pi -\alpha)=\sin (2\pi)\cos \alpha -\cos (2\pi)\sin \alpha \]No w simply simplify this!

OpenStudy (anonymous):

How can you use 2π where alpha goes? Then it's all alphas

OpenStudy (calculusfunctions):

What?? A is 2 pi and B is alpha. You're taking it way too literally! Just simplify and you'll see.

OpenStudy (anonymous):

When this equation was covered in class we were using it for exact value, so we found alpha and beta first as two parts of the original angle (ex: cos(35+45) ) and from there found them as fractions. Then simplify.

OpenStudy (anonymous):

Do I need to do something similar here and find alpha as a fraction to simplify? (This problem was not covered in class..)

OpenStudy (calculusfunctions):

No. Tell me this\[\sin (2\pi)=?\]and\[\cos (2\pi)=?\]

OpenStudy (anonymous):

oh haha sin(2π)= 0 cos(2π)= 1

OpenStudy (calculusfunctions):

Yes, do you understand now?

OpenStudy (anonymous):

\[=(0)(\cos \alpha)-(1)(\sin \alpha)\]

OpenStudy (anonymous):

\[=(\cos \alpha)-(1)(\sin \alpha)\]

OpenStudy (calculusfunctions):

Excuse me?? (0)(cos α) = cos α ????? That would imply that (0)(2) = 2, for example. Did someone change the rules while I was sleeping? LOL

OpenStudy (anonymous):

Woops haha

OpenStudy (calculusfunctions):

OK so you understand?

OpenStudy (anonymous):

Yeah I get it, was just having one of those "no brain" moments lol

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (calculusfunctions):

No problem.

OpenStudy (calculusfunctions):

Thanks for the medal although I'm not entirely sure I deserve that one since I was half a sleep. LOL

OpenStudy (anonymous):

Still better than my teacher's lecture lol

OpenStudy (calculusfunctions):

Thanks. let me know if you need anything else. I'll be online for another half of an hour.

OpenStudy (anonymous):

Alight sounds good, thanks